# Covariant Fractons and Weitzenböck Torsion

### v1

@@ 6 Embedding covariant fractons in linearised MHS theory
where the elements in $\mathscr{M}_{\text{fr}}$ are identified by the invariant fracton field strength because of the gauge invariance of the equations of motion: two solutions $h^{\text{(fr)}}_{\mu\nu}(x)$ and $\tilde{h}^{\text{(fr)}}_{\mu\nu}(x)$ which differ by a fracton gauge transformation are identified in $\mathscr{M}_{\text{fr}}$. Similarly, the moduli space of the linearised mhs theory is the submanifold $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$ of the field space identified by the equations of motion (5.5)–(5.6):
$\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}=\{\mathring{T}_{\mu\nu\varrho}(x)\,:\,\mathring{E}^{\text{(sym)}}_{\mu\nu}(\mathring{T})=0,\;\mathring{E}^{\text{(anti)}}_{\mu\nu}(\mathring{T})=0\}.$ (6.2)
Consider an arbitrary fixed element $f_{\mu\nu\varrho}^{\text{(fr)}}$ in $\mathscr{M}_{\text{lin}}$. If we consider a linearised torsion $\tilde{T}_{\mu\nu\varrho}$, such that
$\tilde{T}_{\mu\nu\varrho}=\tfrac{1}{2}\,f_{\mu\varrho\nu}^{\text{(fr)}},$ (6.3)
which means, in particular, that we are considering a linearised torsion with vanishing totally antisymmetric part, then $\tilde{T}_{\mu\nu\varrho}$ is an element in $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$. To see why, consider that
$\ignorespaces\mathring{E}_{\mu\nu}^{\text{(sym)}}(\tilde{T})$ $=\alpha_{3}\,(R^{(\text{fr})}_{\mu\nu}-\tfrac{1}{2}\,R^{(\text{fr})}\,\eta_{\mu\nu})+\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,f^{(\text{fr})\,\alpha}{}_{(\mu\nu)},$ (6.4a)
$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})$ $=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\partial_{\alpha}\,f^{(\text{fr})\,\alpha}{}_{[\mu\nu]},$ (6.4b)
@@
where $R^{(\text{fr})}_{\mu\nu}$ and $R^{(\text{fr})}$ are given by (2.7) and (2.8). Notice that the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$, which means that, had we varied the action with, say, $\alpha_{2}=0$, we would have obtained the same result. This is not surprising, since there are only two independent quadratic scalar contractions if the totally antisymmetric part of the torsion is vanishing.
The projected symmetric equation (6.4a) corresponds to the equations of motion in covariant fracton gauge theory (2.9)
$\mathring{E}_{\mu\nu}^{\text{(sym)}}(\tilde{T})=E_{\mu\nu}^{\text{(fr)}}(f^{\text{(fr)}})=0,$ (6.5)
upon renaming the constant coefficients
$\alpha=\tfrac{1}{4}\,(2\,\alpha_{1}+\alpha_{2}),\quad\beta=\tfrac{\alpha_{3}}{2}.$ (6.6)
Nevertheless, in linearised mhs theory there is also the variation with respect to the antisymmetric part of the linearised vielbein, which is responsible for the antisymmetric part of the equations of motion (6.4b). But, excluding the case (4.2), which brings to linearised General Relativity and which is such that $02\,\alpha_{1}+\alpha_{2}+\alpha_{3}=$ 0, (6.4b) is implied by (6.4a). To show why, consider the first divergence the fracton equations of motion (2.12). If we define the following one-form Blasi:2022mbl
$j=j_{\mu}\,\mathrm{d}x^{\mu}=\partial^{\nu}\,h^{\text{(fr)}}_{\nu\mu}\,\mathrm{d}x^{\mu},$ (6.7)
the equation (2.12) can be written in this form
$\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}\,\mathrm{d}x^{\nu}=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j=0,$ (6.8)
where $\mathrm{d}$ and $\mathrm{d}^{\dagger}$ are the De Rham differential and codifferential, as it follows by explicit evaluation
$\partial^{\mu}\,\partial^{\varrho}\,f^{\text{(fr)}}_{\varrho(\mu\nu)}\,\mathrm{d}x^{\nu}$ $=2\,\partial^{\mu}\,\partial^{\varrho}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\nu}-\partial^{\mu}\,\partial^{\varrho}\,\partial_{\mu}\,h^{\text{(fr)}}_{\varrho\nu}\,\mathrm{d}x^{\nu}-\partial^{\mu}\,\partial^{\varrho}\,\partial_{\nu}\,h^{\text{(fr)}}_{\varrho\mu}\,\mathrm{d}x^{\nu}=$
$=\,\partial^{2}\,\partial^{\mu}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\nu}-\mathrm{d}x^{\nu}\,\partial_{\nu}\,\partial^{\mu}\,\partial^{\varrho}\,h^{\text{(fr)}}_{\mu\varrho}=$
$=\partial^{2}\,j-\mathrm{d}\,\partial_{\mu}\,j^{\mu}=-(\mathrm{d}^{\dagger}\,\mathrm{d}+\mathrm{d}\,\mathrm{d}^{\dagger})\,j+\mathrm{d}\,\mathrm{d}^{\dagger}\,j=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j.$ (6.9)
Now, the equation (6.8) is equivalent to say that $j$ is $\mathrm{d}$-closed,
$\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}\,\mathrm{d}x^{\nu}=-\mathrm{d}^{\dagger}\,\mathrm{d}\,j=0\Leftrightarrow\mathrm{d}\,j=0,$ (6.10)
[…]
$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}=-\tfrac{1}{2}\,(2\,\alpha_{1}+\alpha_{2}+\alpha_{3})\,\mathrm{d}\,j,$ (6.11)
since
$\partial^{\varrho}\,f^{\text{(fr)}}_{\varrho\mu\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ $=\partial^{\varrho}\,\partial_{\varrho}\,h^{\text{(fr)}}_{\mu\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}-\partial^{\varrho}\,\partial_{\mu}\,h^{\text{(fr)}}_{\varrho\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}=-\mathrm{d}\,j.$ (6.12)
In conclusion, we have shown that
$\mathring{E}_{\mu\nu}^{\text{(anti)}}(\tilde{T})=0\Leftrightarrow\partial^{\mu}\,E_{\mu\nu}^{\text{(fr)}}(f^{(\text{fr})})=0,$ (6.13)
which means that the projected symmetric equation (6.4a) implies the antisymmetric one (6.4b). As a consequence, all the solutions of the covariant fracton gauge theory are solutions of the linearised mhs theory:
@@
$\mathscr{M}_{\text{fr}}\subset\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}.$ (6.14)

### v3

@@ 6 Particle spectrum in covariant fracton gauge theory
This means that two solutions $h^{\text{(fr)}}_{\mu\nu}(x)$ and $h^{\prime\text{(fr)}}_{\mu\nu}(x)$ of the equations of motion, which differ by a fracton gauge transformation are identified in the moduli space.
When $\frac{\beta}{\alpha}=-2$, as in (2.11), the theory reduced to linearised Einstein gravity, and the gauge invariance is extended to all linearised diffeomorphisms:
$\delta_{\text{diff}}\,h_{\mu\nu}^{\text{(fr)}}=\partial_{\mu}\,\xi_{\nu}+\partial_{\nu}\,\xi_{\mu},\;\;\text{if}\;\;\tfrac{\beta}{\alpha}=-2.$ (6.2)
There is also a case in which the fracton theory is traceless, that is, it does not depend on $h^{\text{(fr)}\varrho}{}_{\varrho}$, when $h_{\mu\nu}^{\text{(fr)}}$ is decomposed in its traceless part and in its trace. One can check that this happens if and only if
$\tfrac{\beta}{\alpha}=-\tfrac{2}{d-1},$ (6.3)
where $\alpha$ and $\beta$ are the parameters introduced in (2). In that case, there is an additional gauge symmetry, whose effect is to shift the trace, corresponding to a linearised Weyl scaling of $h^{\text{(fr)}}_{\mu\nu}$:
$\delta_{\text{Weyl}}\,h^{\text{(fr)}}_{\mu\nu}=2\,\sigma\,\eta_{\mu\nu},\;\;\text{if}\;\;\tfrac{\beta}{\alpha}=-\tfrac{2}{d-1}.$ (6.4)
Let us find the general solution of covariant fracton gauge theory. For definiteness, we consider the four dimensional case from now on. The problem was addressed in Afxonidis:2023pdq ; Afxonidis:2024tph , where the particle content of the theory was studied: except for some special cases, the theory describes five propagating degrees of freedom, with helicities $0,\pm 1,\pm 2$ respectively. Working in Fourier space, $p^{\mu}$ denotes the momentum and $\tilde{h}^{\text{(fr)}}_{\mu\nu}$ the Fourier transform of the $h^{\text{(fr)}}_{\mu\nu}$. The equations (2.10) in Fourier space are:
$K_{\mu\nu}{}^{\varrho\sigma}(p^{2};\alpha,\beta)\,\tilde{h}^{\text{(fr)}}_{\varrho\sigma}$ $=2\,\alpha\,p^{2}\,\tilde{h}^{\text{(fr)}}_{\mu\nu}-\tfrac{1}{2}\,(2\,\alpha-\beta)\,p_{\alpha}\,p_{(\mu}\,\tilde{h}^{\text{(fr)}}_{\nu)}{}^{\alpha}\,$+
$-\beta\,p_{\mu}\,p_{\nu}\,\tilde{h}^{\text{(fr)}}+\beta\,\eta_{\mu\nu}\,(p^{2}\,\tilde{h}^{\text{(fr)}}-p_{\alpha}\,p_{\beta}\,\tilde{h}^{\text{(fr)}\,\alpha\beta})=0,$ (6.5)
where $K(p^{2};\alpha,\beta)$, viewed as a $10\times 10$ matrix, identifies a system of homogeneous linear equations in the $10$ variables $\{\tilde{h}^{\text{(fr)}}_{\mu\nu}\}$. Trivial or pure-gauge solutions are those which solve the system without any condition on the momentum, and they correspond to the gauge invariance of the theory. The number of independent trivial solutions is given by the dimension of the kernel of off-shell $K(p^{2};\alpha,\beta)$ (null eigenspace, whose dimension is the number of variables decreased by the matrix rank). One can show that
$\text{trivial solutions}=\dim\,\text{ker}\,K(p^{2}\neq 0;\alpha,\beta)=\begin{cases}6&\text{if}\;\alpha=0,\\ 4&\text{if}\;\tfrac{\beta}{\alpha}=-2,\\ 2&\text{if}\;\tfrac{\beta}{\alpha}=-\tfrac{2}{3},\\ 1&\text{otherwise}.\end{cases}$ (6.6)
$p_{\mu}\,p_{\nu}\,\lambda$ is always a trivial solution; in the linearised Einstein case $\frac{\beta}{\alpha}=-2$, this solution is encompassed in $p_{(\mu}\,\xi_{\nu)}$, which are four independent trivial solutions; in the traceless case $\frac{\beta}{\alpha}=-\frac{2}{3}$, besides $p_{\mu}\,p_{\nu}\,\lambda$, $2\,\eta_{\mu\nu}\,\sigma$ is trivial too. There is one peculiar case more, $\alpha=0$, in which there is no kinetic term for the fracton gauge field, so that the theory is expected to have no propagating degrees of freedom in this case Blasi:2022mbl .
In the massless case, the momentum can be canonically parametrised as $p^{\mu}=(1,0,0,1)$, such that $p^{2}=0$. The dimension of the kernel in the massless case gives the number of independent solutions without taking into account the gauge redundancy. The result is
$\dim\,\text{ker}\,K(p^{2}=0,\alpha,\beta)=\begin{cases}8\quad\text{if}\;\frac{\beta}{\alpha}=2,\\ 6\quad\text{otherwise}.\end{cases}$ (6.7)
The case $\frac{\beta}{\alpha}=2$ is special, because the contraction $p^{\alpha}\,\tilde{h}^{(\text{fr})}_{\alpha\mu}$ disappears from the equations (6.5). Subtracting the number of trivial solutions to the dimension of the kernel in the various cases, one gets the number of degrees of freedom of the theory:
$_{\text{dof}}=\begin{cases}6-6=0,&\text{if}\;\alpha=0,\\ 6-4=2,&\text{if}\;\tfrac{\beta}{\alpha}=-2\;\text{(linearised gravity)},\\ 6-2=4,&\text{if}\;\tfrac{\beta}{\alpha}=-\tfrac{2}{3}\;\text{(traceless limit)},\\ 8-1=7,&\text{if}\;\tfrac{\beta}{\alpha}=2,\\ 6-1=5,&\text{otherwise,}\end{cases}$ (6.8)
where we see that, as expected, the theory has no propagating degrees of freedom when $\alpha=0$. In the case of linearised Einstein gravity, the two degrees of freedom must describe the graviton, with helicities $\pm 2$. In the third case, we expect that the tracelessness condition eliminates a scalar particle, so that in the general case, which has five degrees of freedom, at least a scalar particle is included in the spectrum. Since the general case encompasses also the second, the graviton is also included. Only two degrees of freedom remain to be identified: they turn out to describe a particle with helicities $\pm 1$, as confirmed by the explicit solutions of the equations of motion. The two degrees of freedom more in the special case $\tfrac{\beta}{\alpha}=2$ have again helicities $\pm 1$, as we will see in the end of this section.
Solving the equations of motion means finding a basis for the kernel in the massless case, eliminating the trivial solutions. In the general case, apart from the trivial solutions $p_{\mu}\,p_{\nu}\,\lambda$, the other five independent elements in the basis can be chosen to be equal to
$\text{diag}\,(-1,-\tfrac{2\,\alpha+\beta}{2\,\beta},-\tfrac{2\,\alpha+\beta}{2\,\beta},1),\;\;\lambda_{\mu\nu}^{\pm 1},\;\;\lambda_{\mu\nu}^{\pm 2},\;\;\text{if}\;\beta\neq 0,\alpha\neq 0,\tfrac{\beta}{\alpha}=\pm 2,-\tfrac{2}{3},$ (6.9)
where $\lambda_{\mu\nu}^{s}$ are polarisation matrices with helicity $s=\pm 1,\pm 2$ (the explicit expressions are in (D.4b)–(D.4c)), and the diagonal matrix has vanishing helicity. Therefore, in the general case, covariant fracton gauge theory describes particles with helicities $0,\pm 1,\pm 2$ Afxonidis:2023pdq . If $\beta=0$, the scalar solution in (6.9) is replaced by
$\text{diag}\,(0,1,1,0),$ (6.10)
which again has vanishing helicity. If $\frac{\beta}{\alpha}=-2$, the trivial solutions $p_{(\mu}\,\xi_{\nu)}$ are completed to a basis for the kernel by $\lambda_{\mu\nu}^{\pm 2}$, as expected in linearised gravity. If $\frac{\beta}{\alpha}=-\frac{2}{3}$, the scalar solution in (6.9) becomes $\eta_{\mu\nu}$, since $\frac{2\,\alpha+\beta}{2\,\beta}\rightarrow 1$, so it is trivial, and the theory describes only the modes $\pm 1,\pm 2$. Finally, in the special case $\frac{\beta}{\alpha}=2$, the basis has eight independent elements: one of them is the trivial one $p_{\mu}\,p_{\nu}\,\lambda$; the remaining seven ones can be chosen as
$\text{diag}\,(-1,-1,-1,1),\;\;\begin{pmatrix}0&-1&\mp i&0\\ -1&0&0&0\\ \mp i&0&0&0\\ 0&0&0&0\end{pmatrix},\;\;\begin{pmatrix}0&0&0&0\\ 0&0&0&-1\\ 0&0&0&\mp i\\ 0&-1&\mp i&0\end{pmatrix},\;\;\lambda_{\mu\nu}^{\pm 2},$ (6.11)
where the diagonal solution has vanishing helicity, $\lambda_{\mu\nu}^{\pm 2}$ has helicities $\pm 2$, and the remaining two pairs of solutions have both helicities $\pm 1$. So, compared to the general case, there is one $\pm 1$ helicity particle more.
As a last observation, consider that, the polarisation matrices $\lambda_{\mu\nu}^{s}$ (D.4a)–(D.4c), in the general solution (6.9), are transversal in the massless case, meaning that
$p^{\mu}\,\lambda_{\mu\nu}^{s}=0,\;\;s=\pm 1,\pm 2,\;\;\text{if}\;p^{\mu}=(1,0,0,1).$ (6.12)
Instead, the scalar mode is longitudinal
$p^{\mu}\,\text{diag}\,(-1,-\tfrac{2\,\alpha+\beta}{2\,\beta},-\tfrac{2\,\alpha+\beta}{2\,\beta},1)_{\mu\nu}=p_{\nu}.$ (6.13)
Nonetheless, this implies that the arbitrary solution $\tilde{h}_{\mu\nu}^{(\text{fr})}$ on-shell satisfies the condition
$p^{\alpha}\,p_{[\mu}\,\tilde{h}^{\text{(fr)}}_{\nu]\alpha}=0\;\;\text{if}\;p^{\mu}=(1,0,0,1),$ (6.14)
which will be useful in the following.
7 Particle spectrum of MHS theory and fracton embedding
Let us now find the particle spectrum for the mhs theory (see also Hayashi:1979qx ), following the same method as in the previous section. It will be convenient to redefine the constants $\alpha_{1}$ and $\alpha_{2}$ as
$\tilde{\alpha}_{1}=2\,\alpha_{1}+\alpha_{2},\quad\tilde{\alpha}_{2}=2\,\alpha_{1}-\alpha_{2}.$ (7.1)
The moduli space $\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}$ of the linearised mhs theory is the submanifold in the field space $\{h_{\mu\nu}$, $b_{\mu\nu}\}$ identified by the equations of motion (5.5)–(5.6), modulo the gauge invariance, which amounts to linearised diffeomorphisms acting both on $h_{\mu\nu}$ and $b_{\mu\nu}$, [^14] In this section $\xi_{\mu}$, and $\Omega_{\mu\nu}$ are simple gauge parameters. We maintain the same notation of the corresponding ghosts (omitting the dot denoting the linearisation) because this should not be cause confusion.
$\delta_{\text{diff}}\,h_{\mu\nu}=\partial_{\mu}\,\xi_{\nu}+\partial_{\nu}\,\xi_{\mu},\;\delta_{\text{diff}}\,b_{\mu\nu}=\partial_{\mu}\,\xi_{\nu}-\partial_{\nu}\,\xi_{\mu},\;\;\forall\;\tilde{\alpha}_{1},\tilde{\alpha}_{2},\alpha_{3},$ (7.2)
as in (5.15). As in covariant fracton gauge theory, there are some particular cases in which the gauge symmetry is enlarged. Linearised Einstein gravity is reproduced when the parameters are as in (4.2), or equivalently $\tilde{\alpha}_{2}=0,\frac{\alpha_{3}}{\tilde{\alpha}_{1}}=-1$, according to (7.1), up to an overall constant. In this case there is no dependence on $b_{\mu\nu}$, so it can be shifted away, and this is the effect of local Lorentz invariance, as previously discussed:
$\delta_{\text{Lorentz}}\,h_{\mu\nu}=0,\;\delta_{\text{Lorentz}}\,b_{\mu\nu}=\Omega_{\mu\nu},\;\;\text{if}\;\;\tilde{\alpha}_{2}=0,\;\tfrac{\alpha_{3}}{\tilde{\alpha}_{1}}=-1.$ (7.3)
The traceless limit in mhs theory is reached if and only if
$\tfrac{\alpha_{3}}{\tilde{\alpha}_{1}}=-\tfrac{1}{d-1}.$ (7.4)
in which case there is no dependence on the trace $h^{\varrho}{}_{\varrho}$ and the gauge invariance includes a Weyl scaling:
$\delta_{\text{Weyl}}\,h_{\mu\nu}=2\,\sigma\,\eta_{\mu\nu},\;\;\delta_{\text{Weyl}}\,b_{\mu\nu}=0,\;\;\text{if}\;\;\tfrac{\alpha_{3}}{\tilde{\alpha}_{1}}=-\tfrac{1}{d-1}.$ (7.5)
A last limit case is when the theory does not depend on $h_{\mu\nu}$ at all. This happens when $\tilde{\alpha}_{1}=\alpha_{3}=0$, in which case the symmetric equation (5.5) trivialises, and mhs theory reduces to the theory for a free two-form gauge field (Ramond-Kalb field rk Kalb:1974yc ):
$\delta_{\textsc{rk}}\,h_{\mu\nu}=\Delta_{\mu\nu},\;\delta_{\textsc{rk}}\,b_{\mu\nu}=\partial_{\mu}\,\xi_{\nu}-\partial_{\nu}\,\xi_{\mu},\;\;\text{if}\;\;\tilde{\alpha}_{1}=\alpha_{3}=0,$ (7.6)
where $\Delta_{\mu\nu}$ is any symmetric rank-two tensor, shifting $h_{\mu\nu}$.
In order to find the general solution of (5.5)–(5.6), we write them in Fourier space. Again, we consider from now on the four-dimensional case. The equations read
$\tfrac{1}{2}\,\tilde{\alpha}_{1}\,p^{2}\,\tilde{h}_{\mu\nu}-\tfrac{1}{4}\,(\tilde{\alpha}_{1}-\alpha_{3})\,p_{\alpha}\,p_{(\mu}\,\tilde{h}_{\nu)}{}^{\alpha}-\tfrac{1}{2}\,\alpha_{3}\,p_{\mu}\,p_{\nu}\,\tilde{h}\,$+
$\quad+\tfrac{1}{2}\,\alpha_{3}\,\eta_{\mu\nu}\,(p^{2}\,\tilde{h}-p_{\alpha}\,p_{\beta}\,\tilde{h}^{\alpha\beta})+\tfrac{1}{4}\,(\tilde{\alpha}_{1}+\alpha_{3})\,p_{\alpha}\,p_{(\mu}\,\tilde{b}_{\nu)}{}^{\alpha}=0,$ (7.7a)
$\tfrac{1}{2}\,\tilde{\alpha}_{2}\,p^{2}\,\tilde{b}_{\mu\nu}-\tfrac{1}{4}\,(\tilde{\alpha}_{1}-2\,\tilde{\alpha}_{2}+\tilde{\alpha}_{3})\,p^{\alpha}\,p_{[\mu}\,\tilde{b}_{\nu]\alpha}-\tfrac{1}{4}\,(\tilde{\alpha}_{1}+\alpha_{3})\,p^{\alpha}\,p_{[\mu}\,\tilde{h}_{\nu]\alpha}=0.$ (7.7b)
This is a system of homogeneous linear equations in the $10+6$ variables $(\tilde{h}_{\mu\nu},\tilde{b}_{\mu\nu})$.
@@ 7 Particle spectrum of MHS theory and fracton embedding
Looking at the equations alone it is possible to deduce the embedding of the fracton moduli space in the mhs one. Indeed, one sees that, if [^15] In direct space, this means that the linearised torsion $\mathring{T}_{\mu\nu\varrho}$, corresponding to the choice (7.8), $\mathring{T}_{\mu\nu\varrho}=\tfrac{1}{2}\,f_{\mu\varrho\nu}^{\text{(fr)}}$. Moreover, there is no dependence in the projected equations on $\tilde{\alpha}_{2}$, and that the old constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}=\tilde{\alpha}_{1}$, which means that, had we varied the action with, say, $\alpha_{2}=0$, we would have obtained the same result. This is not surprising, since there are only two independent quadratic scalar contractions if the totally antisymmetric part of the torsion is vanishing.
$\tilde{h}_{\mu\nu}=\tilde{h}_{\mu\nu}^{\text{(fr)}},\quad\tilde{b}_{\mu\nu}=0,\quad\tilde{\alpha}_{1}=4\,\alpha,\quad\alpha_{3}=2\,\beta,$ (7.8)
then the first equation (7.7a) reduces to the fracton equation (6.5), and the second equation (7.7b) reduces to the condition (6.14). This shows that the fracton moduli space, if $\frac{\beta}{\alpha}\neq 2$, is contained in the subsector with vanishing $b_{\mu\nu}$ of the linearised mhs moduli space, upon using the identification in (7.8),
@@
$\mathscr{M}_{\text{fr}}(\tfrac{\beta}{\alpha}\neq 2)\subseteq{\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}}_{\big|_{b=0}}\subset\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}.$ (7.9)
Observe that the condition $b_{\mu\nu}=0$ in (7.8) breaks the gauge invariance (7.2), but it leaves a residual gauge freedom when $\xi_{\mu}$ is exact, that is, when it generates longitudinal diffeomorphisms $\xi_{\mu}\propto\partial_{\mu}\,\lambda$. This is consistently the fracton gauge invariance (6.1).
Actually, we can prove an even stronger statement: not only the fracton moduli space is included in the $b_{\mu\nu}=0$ subsector, but it is also isomorphic to the same subsector. To see why, let us find the most general solution of the full equations (7.7a)–(7.7b). At the very end, we will see that, upon restricting in the $b_{\mu\nu}=0$ subsector, the most general solutions coincides with the fracton one, if $\frac{\beta}{\alpha}\neq 2$.
A $16\times 16$ matrix $K(p;\tilde{\alpha}_{1},\tilde{\alpha}_{2},\alpha_{3})$ can be introduced, describing the system (7.7a)–(7.7b) in the variables $(\tilde{h}_{\mu\nu},\tilde{b}_{\mu\nu})$, viewed as a vector with $10+6$ entries. Except for the special cases in which the gauge invariance is enlarged, the dimension of the kernel in the massless case is 10, and the number of trivial solutions is 4, encoded in $(\tilde{h}_{\mu\nu},\tilde{b}_{\mu\nu})=(p_{(\mu}\,\xi_{\nu)},p_{[\mu}\,\xi_{\nu]})$. So, there are six degrees of freedom, which can be chosen to be equal to
$(\text{diag}\,(-1,-\tfrac{\tilde{\alpha}_{1}+\alpha_{3}}{2\,\alpha_{3}},-\tfrac{\tilde{\alpha}_{1}+\alpha_{3}}{2\,\alpha_{3}},1),O),\;\;(\lambda_{\mu\nu}^{\pm 1},O),\;\;(\lambda_{\mu\nu}^{\pm 2},O),\;\;(O,\lambda_{\mu\nu}^{0}),$ (7.10)
where $O$ is the $4\times 4$ null matrix, and $\lambda_{\mu\nu}^{s}$ are the polarisation matrices in (D.4a)–(D.4c), with helicities $s=0,\pm 1,\pm 2$. Therefore, mhs theory describes particles with helicities $0,\pm 1,\pm 2$, encoded in the symmetric part $\tilde{h}_{\mu\nu}$, and a scalar particle, encoded in the antisymmetric part $\tilde{b}_{\mu\nu}$ – this is the expected result, since, as well known, a two-form gauge field in four dimensions is dual to a scalar particle. Moreover, the subsector with vanishing antisymmetric part is the same as in covariant fracton gauge theory, when the special case $\frac{\beta}{\alpha}=2$ is excluded, since, using the identification of the constants in (7.8), $\frac{\tilde{\alpha}_{1}+\alpha_{3}}{2\alpha_{3}}=\frac{2\alpha+\beta}{2\beta}$, so that the first solution is the same as the first solution in (6.9). Therefore, the statement in (7.9) holds also in the stronger version
$\mathscr{M}_{\text{fr}}(\tfrac{\beta}{\alpha}\neq 2)={\mathscr{M}_{\textsc{mhs}}^{\text{(lin)}}}_{\big|_{b=0}}.$ (7.11)